# Factorizable joint shift revisited

### v1

@@ 3.2 Using FJS for inference about the target distribution
    In addition to the assumptions of (i), suppose that the source distribution $P$ on $(\Omega,\mathcal{F})$ satisfies Assumptions 2.4 (i) and (ii) and define $\psi(y)=\frac{\overline{g}(y)}{g(y)}$ for $y\in\Omega_{Y}$. Then (3.5a) is equivalent to
@@
    This observation implies $P\bigl[\overline{g}(Y)>0,E_{P}[\overline{g}(Y)\,|\,\sigma(X)=0]=0\bigr]=0$. Taking into account $P[\overline{g}(Y)>0]=1$, it follows that $P\bigl[E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\bigr]=0$. Therefore, division by $E_{P}[\overline{g}(Y)\,|\,\sigma(X)]$ is possible and the equivalence of (3.3) and (3.5a) follows.
@@
  • Transformed into the notation of Corollary 3.7, (9a) of Tasche (2022) reads $b=\overline{g}(Y)$, with $\frac{Q[Y=y]}{P[Y=y]}=g(y)$ and $\varrho(y)=\frac{\psi(y)}{\psi(d)}$ for $y=1,\ldots,d$.
@@
In the setting of Section 2.1, suppose that Assumptions 2.2 and 2.4 (ii) are satisfied and that $Q$ and $P$ are related through FJS with $\frac{dQ}{dP}=\overline{h}(X)\,\overline{g}(Y)>0$ $P$-a.s.
@@
with $\psi=\frac{\overline{g}}{g}$ as in Corollary 3.7 (ii). By (3.6), the ratio of the target conditional feature densities evaluated at the same feature $x$ but for two different labels $y_{0}$, $y_{1}$ does not depend on the value of the feature. In other words, $q_{X|Y=y_{1}}$ and $q_{X|Y=y_{0}}$ differ only by a calibration factor and are even identical in the case $\overline{g}=g$ (or $\psi=1$).
@@
With the notation of Corollary 3.7 (ii) and assuming $\frac{dQ_{Y}}{dP_{Y}}=g>0$, Corollary 2.5 (ii) implies the posterior correction formula for general label spaces under FJS:
@@
(3.7) also follows from combining (2.6c) with (3.6). Eq. (12) of Tasche (2022) is the special case of (3.7) with $G=\{j\}$ for categorical label spaces.
@@
Under the assumption of FJS, Corollary 3.7 also suggests an ansatz for solving Problem (B) as phrased in Section 2. Recall that in Problem (B) only the feature marginal distribution $Q_{X}$ (or its density $h$ with respect to $P_{X}$) is known of the source distribution $Q$.
@@
  • Solve the integral equation (3.5c) with $\psi=1$ for the second marginal density $g$ under the constraints $g\geq 0$ and $\int g(x)\,P_{X}(dx)=1$ to obtain the density $g$ of $Q_{Y}$ with respect to $P_{Y}$. Again, it might be preferable to make use of the simplied equation (3.8).
@@
(3.4a) together with Corollary 2.5 (ii) also could be the basis for a ‘general label space’ variant of the expectation-maximisation (EM) algorithm for estimating target prior class probabilities (Saerens et al., 2002, Eq. (2.9)). This ansatz for determining the target label density $g$ of $Q_{Y}$ with respect to $P_{Y}$ reads as follows:
@@
$\begin{split}q_{Y|X=x}^{(n)}(y)&=\ \frac{g_{n}(y)}{\int g_{n}(z)\,\varphi(x,z)\,P_{Y}(dz)},\quad y\in\Omega_{Y},x\in\Omega_{X},\\ g_{n+1}(y)&=\ \int q_{Y|X=x}^{(n)}(y)\,\varphi(x,y)\,Q_{X}(dx),\quad y\in\Omega_{Y}.\end{split}$ (3.11)
@@ 4 Generalized label shift (GLS)
$0\ =\ Q\bigl[E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\bigr]\ =\ E_{P}\bigl[E_{P}[f(X,Y)\,|\,\sigma(X)]\,\mathbf{1}_{E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0}\bigr].$ (4.3b)
@@
But then by Corollary 4.2, if $Q$ and $P$ are related through GLS they also are related through FJS, even with the same density of the features conditional on the label for all label values as discussed in Remark 3.9. This observation applies independently of the representation mapping $R$. Hence it might make sense to renounce finding a representation $R$ and instead resort to the ansatz for solving Problem (B) as presented at the end of Section 3.2.
5 Conclusions
@@
In order to facilitate the related analyses, we have introduced an assumption on the source distribution with regard to the feasibility of regular conditional distributions of the labels given the features and vice versa. Another assumption on the existence of a density of the source distribution with respect to the product measure of its feature and label marginal distributions has turned out to be useful for the reconciliation of the results with results for categorical labels as well as for calculations in practice.
With regard to fitting FJS to given target feature and label marginal distributions, we have presented an approach based on a non-linear integral equation for which we have proposed an iterative solution. Regarding fitting FJS to a given target feature distribution without recourse to any label observation, to cover the case of general label spaces we have proposed a generalisation of the expectation maximisation (EM) approach for estimating target prior class probabilities (Saerens et al., 2002).
Future research topics could be the computational properties of the estimation approaches proposed in Section 3.2 and the extension of other specific types of distribution shift (e.g. sparse joint shift, Chen et al., 2022) to include the case of general label spaces.

### v2

@@ 3.2 Solving Problems (A) and (B) by assuming FJS
    In addition to the assumptions of (i), suppose that the source distribution $P$ on $(\Omega,\mathcal{F})$ satisfies Assumptions 2.4 (i) and (ii) and define $\psi(y)=\frac{\overline{g}(y)}{g(y)}$ for $y\in\Omega_{Y}$. Then (3.5a) and (3.3) are equivalent to
@@
    This observation implies $P\bigl[\overline{g}(Y)>0,E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\bigr]=0$. Taking into account $P[\overline{g}(Y)>0]=1$, it follows that $P\bigl[E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\bigr]=0$. Therefore, division by $E_{P}[\overline{g}(Y)\,|\,\sigma(X)]$ is possible and the equivalence of (3.3) and (3.5a) follows.
@@
  • Transformed into the notation of Corollary 3.7, (9a) of Tasche (2022) reads $b=\overline{g}(Y)$, with $\frac{Q[Y=y]}{P[Y=y]}=g(y)$ and $\varrho(y)=\frac{\psi(y)}{\psi(d)}$ for $y=1,\ldots,d-1$.
@@
In the setting of Section 2.1, suppose that Assumption 2.4 (ii) is satisfied and that $Q$ and $P$ are related through FJS with $\frac{dQ}{dP}=\overline{h}(X)\,\overline{g}(Y)>0$ $P$-a.s.
@@
with $\psi=\frac{\overline{g}}{g}$ as in Corollary 3.7 (ii). By (3.6), the ratio of the target conditional feature densities evaluated at the same feature $x$ but for two different labels $y_{0}$ and $y_{1}$ does not depend on the value of the feature. In other words, $q_{X|Y=y_{1}}$ and $q_{X|Y=y_{0}}$ differ only by a calibration factor and are even identical in the case $\overline{g}=g$ (or $\psi=1$).
@@
With the notation of Corollary 3.7 (ii), under Assumption 2.4 (i) (3.4a) of Remark 3.6 implies the posterior correction formula for general label spaces under FJS:
@@
(3.7) also follows from combining (2.6c) with (3.6). Eq. (12) of Tasche (2022) is the special case of (3.7) for categorical label spaces with $G=\{j\}$.
3.2.1 Solving Problem (A) by assuming FJS
@@
Note that the existence of $\lim_{n\to\infty}\psi_{n}(y)=\psi^{\ast}(y)$ for all $y$ of the sequence defined by (3.9) does not necessarily always imply that $\psi^{\ast}$ fulfills (3.8) because without further conditions it cannot be excluded that
$\lim_{n\to\infty}\int\frac{\varphi(x,y)}{\int\psi_{n}(z)\,\varphi(x,z)\,Q_{Y}(dz)}\,Q_{X}(dx)\ \neq\ \int\frac{\varphi(x,y)}{\int\lim_{n\to\infty}\psi_{n}(z)\,\varphi(x,z)\,Q_{Y}(dz)}\,Q_{X}(dx).$ (3.10)
Clearly, (3.10) is not a problem in the case of a categorical label space. Tasche (2022, Proposition 2) proved existence and uniqueness of a solution to (3.8) for the case of binary labels.
3.2.2 Solving Problem (B) by assuming covariate shift
A straightforward way to solving Problem (B) as formulated in Section 2 is by assuming that the target distribution $Q$ and the source distribution $P$ are related through covariate shift in the sense of Definition 3.2. Covariate shift is implied by “missing at random” sample selection bias (Moreno-Torres et al., 2012, Section 6.1) such that there are applications where covariate shift is the only reasonable assumption.
Problem (B) means that a target feature distribution $Q^{\ast}_{X}$ is pre-specified by its density $h$ with respect to the source feature distribution $P_{X}$. According to Proposition 3.3, then letting $\frac{dQ}{dP}=h(X)$ defines a joint distribution $Q$ of features $X$ and labels $Y$ such that $Q$ and $P$ are related through covariate shift and it holds that $Q_{X}=Q^{\ast}_{X}$. Thus, Problem (B) is solved in principle.
However, Storkey (2009, Section 5.1) commented that $\ldots$ there is some benefit to be obtained by doing some-thing different in the case of covariate shift. The argument here is that these papers indicate a computational benefit rather than a fundamental modelling benefit.” The computational aspects of covariate shift have been extensively studied in the literature and are not discussed in this paper. See for example Kouw and Loog (2021, Section 3.1) for a review of the literature on domain adaptation under covariate shift.
3.2.3 Solving Problem (B) by assuming label shift or constrained FJS
Recall that in Problem (B) as phrased in Section 2 only the feature marginal distribution $Q_{X}$ (or its density $h$ with respect to $P_{X}$) is known of the source distribution $Q$. While most of the time the aim in Problem (B) is to come up with an estimate of the target posterior distribution $Q_{Y|X}$ of the labels conditional on the features, thanks to the law of total probability the marginal label distribution $Q_{Y}$ then is a byproduct. In the literature, the quest for $Q_{Y}$ in the case of categorical labels is called ‘quantification’ ‘counting’, ‘class probability re-estimation’, ‘class prior estimation’ or ‘class distribution estimation’ (Esuli et al., 2023, p. 2). In the case of general normed label spaces, Azizzadenesheli (2022) referred to the estimation of $Q_{Y}$ as ‘Importance Weight Estimation’.
Under the assumption of FJS, Corollary 3.7 and Theorem 3.5 (ii) together suggest the following ansatz for solving Problem (B).
@@ 3.2.3 Solving Problem (B) by assuming label shift or constrained FJS
  • Solve the integral equation (3.5c) with $\psi=1$ for the function $g$ under the constraints $g\geq 0$ and $\int g(y)\,P_{Y}(dy)=1$ to obtain the density $g$ of the label distribution $Q_{Y}$ with respect to $P_{Y}$. Again, it might be preferable to make use of the simplied equation (3.8) if possible.
@@
For the case of categorical labels, Tasche (2014, Theorem 1 (b)) provided a result on the uniqueness of the solution $g$ of (3.5c) with $\psi=1$. Peters and Coberly (1976) analysed the properties of a fixed point version of (3.5c) in the case of categorical labels and analysed an iterative solution approach.
(3.4a) together with Corollary 2.5 (ii) also is the basis for a ‘general label space’ variant of the expectation-maximisation (EM) algorithm for estimating target prior class probabilities (Saerens et al., 2002, Eq. (2.9)). This method for determining the target label density $g$ of $Q_{Y}$ with respect to $P_{Y}$ reads as follows:
@@
$\begin{split}q_{Y|X=x}^{(n)}(y)&=\ \frac{g_{n}(y)}{\int g_{n}(z)\,\varphi(x,z)\,P_{Y}(dz)},\quad y\in\Omega_{Y},x\in\Omega_{X},\\ g_{n+1}(y)&=\ \int q_{Y|X=x}^{(n)}(y)\,\varphi(x,y)\,Q_{X}(dx),\quad y\in\Omega_{Y},\end{split}$ (3.12)
with $\varphi$ as defined in (2.8).
In the following analyses of the EM algorithm specified by (3.11a) and (3.11b), we make use of the standard notion of Kullback-Leibler divergence (cf. Section 1.6.1 of Bishop, 2006).
Definition 3.11.
Let $(\overline{\Omega},\mathcal{M},\mu)$ be a probability space. For $\mathcal{M}$-Borel measurable functions $\lambda_{0}\geq 0$ and $\lambda_{1}\geq 0$ with $\int\lambda_{i}\,d\mu=1$, $\mu[\lambda_{i}=0]=0$ for $i=0,1$ and $\int\lambda_{0}\left|\log\left(\frac{\lambda_{0}}{\lambda_{1}}\right)\right|d\mu<\infty$, the Kullback-Leibler divergence $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})$ of $\lambda_{1}$ with respect to $\lambda_{0}$ is defined as $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})=\int\lambda_{0}\,\log\left(\frac{\lambda_{0}}{\lambda_{1}}\right)d\mu$.
Jensen’s inequality implies $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})\geq 0$ and $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})=0$ if and only if $\lambda_{0}=\lambda_{1}$ $\mu$-a.s. Accordingly, $\lambda_{0}$ and $\lambda_{1}$ are the more similar, the closer to naught $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})$ is. Minimising $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})$ by changing $\lambda_{1}$ may be interpreted as a population version of maximum likelihood estimation.
Theorem 3.12 generalises the appendix of Saerens et al. (2002) “Derivation of the EM Algorithm” from categorical label spaces to general label spaces. Thus, for non-categorical label spaces, Theorem 3.12 establishes the EM algorithm as a potential alternative to the ‘inverse operator’ approach (comparable to ‘adjusted classify & count’ in the case of categorical labels) by Azizzadenesheli (2022). Proposition 3.14 generalises Theorem 1 on ‘exact fit’ of Tasche (2014). Proofs for Theorem 3.12 and Proposition 3.14 are presented below in Appendix B.
Theorem 3.12.
Under Assumptions 2.4 (i) and (ii), let $(\Omega,\mathcal{F},P)=(\Omega_{X}\times\Omega_{Y},\sigma(X,Y),P_{(X,Y)})$ be a source distribution of features $X$ and labels $Y$ as defined in Section 2.1. Assume that there is a some feature distribution $Q^{\ast}_{X}$ with a density $h>0$ with respect to $P_{X}$. In addition [^7] A further implicit assumption for Theorem 3.12 is that all integrals used in (3.13) and the proof of the theorem exist and are finite. In particular, the assumptions $E_{P}\bigl[h(X)\,|\log(h(X)|\bigr]<\infty$ and $E_{P}\bigl[h(X)\,|\log(h_{n}(X)|\bigr]<\infty$ for $n=0,1,\ldots$ ought to be made for Theorem 3.12 (ii)., suppose there is an initial label density $g_{0}>0$ with respect to $P_{Y}$. Then the following statements hold true:
  • (i)
    All functions $q^{(n)}_{Y|X=x}$ for $x\in\Omega_{X}$ and $n=0,1,\ldots$ and $g_{n}$ for $n=1,2,\ldots$ defined recursively by (3.11a) and (3.11b) are positive and densities with respect to $P_{Y|X=x}$ and $P_{Y}$ respectively.
  • (ii)
    Define a sequence $Q^{(n)}$, $n=0,1,\ldots$, of target distributions on $(\Omega,\mathcal{F})$ by letting $\frac{dQ^{(n)}}{dP}=g_{n}(Y)$ such that by Proposition 3.3 all $Q^{(n)}$ are related to $P$ through label shift. With $h_{n}=\frac{dQ^{(n)}_{X}}{dP_{X}}=E_{P}[g_{n}(Y)\,|\,X=\cdot]$, then it holds that
    $\mathrm{KL}_{P_{X}}(h\parallel h_{n+1})\ \leq\ \mathrm{KL}_{P_{X}}(h\parallel h_{n}),\quad\text{for all}\ n=0,1,\ldots.$ (3.13)
Remark 3.13 (Interpretation of Theorem 3.12).
Theorem 3.12 describes the EM algorithm as specified by (3.11a) and (3.11b) as the construction of a sequence of target joint distributions $Q^{(n)}$ of features $X$ and labels $Y$ such that $Q^{(n)}$ and the source joint distibution $P$ are related through label shift. Each $Q^{(n)}$ entails a target feature distribution $Q^{(n)}_{X}$ with a $P_{X}$-density $h_{n}$. As a result, there is an accompanying sequence of Kullback Leibler divergences $\mathrm{KL}_{P_{X}}(h\parallel h_{n})$ of $Q^{(n)}_{X}$ with respect to the pre-specified feature distribution $Q^{\ast}_{X}$ where both feature distributions are represented through their respective densities with respect to the source feature distribution $P_{X}$. By (3.13), this sequence decreases such that $Q^{(n+1)}_{X}$ resembles $Q^{\ast}_{X}$ more than $Q^{(n)}_{X}$. Hence stopping the EM algorithm at any step does not result in a deterioration of the approximation of the pre-specified feature distribution $Q^{\ast}_{X}$ when compared to the approximation achieved before.
Although the decreasing sequence $\mathrm{KL}_{P_{X}}(h\parallel h_{n})$ converges to a limit $\mathrm{KL}^{\ast}$ for $n\to\infty$, Theorem 3.12 does not exclude the possibility that $\mathrm{KL}^{\ast}>0$ which would suggest that $Q^{\ast}_{X}$ does not correspond to any target distribution related to the source distribution $P$ through label shift. Neither does Theorem 3.12 state that there is a density $h^{\ast}$ with $\mathrm{KL}_{P_{X}}(h\parallel h^{\ast})=\mathrm{KL}^{\ast}$ or that the $Q^{(n)}$ converge to any limit distribution. For the case of a categorical label space, Wu (1983) provided sufficient conditions for the sequence $g_{n}$, $n=0,1,\ldots$, of target prior class probabilities to converge which would entail convergence of the related target joint distributions $Q^{(n)}$.
For the case of a general label space, the following proposition presents consequences of the potential convergence of the target label densities $g_{n}$ which are constructed with the EM algorithm as described by (3.11a) and (3.11b).
Proposition 3.14.
In addition to the assumptions of Theorem 3.12, suppose that for the densities defined in Theorem 3.12 (i) it holds that $\lim_{n\to\infty}g_{n}(y)=g(y)>0$ exists $P_{Y}$-a.s. for $y\in\Omega_{Y}$ and $\lim_{n\to\infty}q^{(n)}_{Y|X=x}(y)=q_{Y|X=x}(y)>0$ exists $P$-a.s. for $(x,y)\in\Omega$. Assume also [^8] The assumptions on the exchangeability of integration and $\lim$ in particular hold true if there is some $M>0$ such that $g_{n}\leq M$ and $q_{Y|X=x}^{(n)}\leq M$ for all $n=0,1,\ldots$ and $x\in\Omega_{X}$. that $\int g(y)\,P_{Y}(dy)=1$,
$\begin{split}\lim\limits_{n\to\infty}\int g_{n}(z)\,P_{Y|X=x}(dz)&=\int g(z)\,P_{Y|X=x}(dz)\quad P_{X}\text{-a.s.\ for}\ x\in\Omega_{X},\quad\text{and}\\ \lim\limits_{n\to\infty}\int q^{(n)}_{Y|X=x}(y)\,h(x)\,P_{X|Y=y}(dx)&=\int q_{Y|X=x}(y)\,h(x)\,P_{X|Y=y}(dx)\quad P_{Y}\text{-a.s.\ for}\ y\in\Omega_{Y}.\end{split}$ (3.14)
Then the following statements hold true:
  • (i)
    $g$ is a density with respect to $P_{Y}$. The joint distribution $\widehat{Q}$ of $X$ and $Y$ defined by $\frac{d\widehat{Q}}{dP}=g(Y)$ is related to $P$ through label shift such that $\frac{d\widehat{Q}_{Y}}{dP_{Y}}=g$ follows. With $\widehat{h}(x)=\frac{d\widehat{Q}_{X}}{dP_{X}}(x)=E_{P}[g(Y)\,|\,X=x]$ for $x\in\Omega_{X}$ it holds that
    $\mathrm{KL}_{P_{X}}(h\parallel\widehat{h})\leq\ \inf_{n}\mathrm{KL}_{P_{X}}(h\parallel h_{n})$ (3.15)
    for the feature densities $h_{n}=E_{P}[g_{n}(Y)\,|\,X=\cdot]$.
  • (ii)
    $g$ is a solution of (3.5c) in the case $\psi=1$. Define $f(x,y)=\frac{g(y)\,h(x)}{\int g(z)\,P_{Y|X=x}(dz)}$ for $(x,y)\in\Omega$ $P$-a.s. Then $f$ is a density with respect to $P$. The joint distribution $Q$ of $X$ and $Y$ defined by $\frac{dQ}{dP}=f$ is related to $P$ through FJS in the sense of Definition 3.1 and it holds that $Q_{X}=Q_{X}^{\ast}$ and $\frac{dQ_{Y}}{dP_{Y}}=g$.
Proposition 3.14 (ii) confirms for general label spaces that if the sequence of densities $g_{n}$ constructed with the EM algorithm converges to a positive limit density $g$, then this density $g$ solves (3.5c) in the case $\psi=1$ and as a consequence also Problem (B) as formulated in Section 2. Thereby, Proposition 3.14 (ii) also confirms for general label spaces the ‘exact fit’ property proven by Tasche (2014, Remark 1) for the case of categorical labels. ‘Exact fit’ as discussed by Tasche (2014) means that there is a target joint distribution such that the ratios of pairs of feature densities conditional on the labels do not depend on feature values (cf. Remark 3.9 above regarding this property for FJS in the case of non-categorical labels) and that the target marginal feature distribution exactly matches a pre-specified or observed feature distribution.
Note that exact fit in the sense of matching a target feature distribution with a constrained FJS (by requiring $\psi=1$ in (3.5c)) is not always possible as pointed out by Titterington et al. (1985, Example 4.3.1) in the context of maximum likelihood estimation of class prior probabilities in the case of binary labels. This observation does not contradict Proposition 3.14 (ii) because in cases of imperfect fit one or more of the conditions listed in the proposition would be violated.
[…]
Does imperfect fit of the pre-specified feature distribution by an approximate solution of Problem (B) matter in practice?
  • The answer is no if it is the target posterior label distribution given the features that is important. This assessment follows from Remark 3.10 that shows that the target posterior feature distribution given the label is the same for label shift and for more general FJS with $\psi=1$ as long as the target label distribution is the same.
  • The answer is yes if it is the target posterior feature distribution given the label that is important. Under label shift (i.e. when the fit is imperfect) the target and source posterior feature distributions given the label are equal. This need not be the case when there is exact fit thanks to FJS with $\psi=1$ as in Proposition 3.14 (ii). This observation is a consequence of Remark 3.9 which demonstrates that the target and source posterior feature distributions given the label may observably differ. See Figure 1 of Tasche (2017) for an example of this phenomenon.
@@ 4 Generalized label shift (GLS)
$0\ =\ Q\bigl[E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\bigr]\ =\ E_{P}\bigl[E_{P}[f(X,Y)\,|\,\sigma(X)]\,\mathbf{1}_{\{E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\}}\bigr].$ (4.3b)
@@
But then by Corollary 4.2, if $Q$ and $P$ are related through GLS they also are related through FJS, even with the same density of the features conditional on the label for all label values as discussed in Remark 3.9, i.e. $\psi=1$ in (3.5c). This observation applies independently of the representation mapping $R$. Hence it might make sense to renounce finding a representation $R$ for achieving GLS and instead resort to the EM algorithm for solving Problem (B) as presented in Section 3.2.3.
5 Concluding remarks
@@
In order to facilitate the related analyses, we have introduced an assumption on the source distribution with regard to the feasibility of regular conditional distributions of the labels given the features and vice versa. Another assumption on the existence of a density of the source distribution with respect to the product measure of its feature and label marginal distributions has turned out to be useful for the reconciliation of the theoretical results with results for categorical labels as well as for calculations in practice.
With regard to fitting FJS to given target feature and label marginal distributions, we have presented an approach based on a non-linear integral equation for which we have proposed an iterative solution. Regarding fitting FJS to a given target feature distribution without recourse to any label observations, in order to cover the case of a general label space we have presented a generalisation of the expectation maximisation (EM) approach for estimating target prior class probabilities (Saerens et al., 2002) and shown that it is fit for purpose.
Future research topics could be the computational properties of the estimation approaches proposed in Sections 3.2.1 and 3.2.3 and the study of other specific types of distribution shift (e.g. sparse joint shift, Chen et al., 2022) in the face of general label spaces.